OpenStudy (anonymous):

Find the Maclaurin expansion for the interval from 0 to x (e^t - 1)/t dt

OpenStudy (anonymous):

\[\int\limits_{0}^{x} (e^t - 1)/t dt\]

OpenStudy (anonymous):

first write the expansion for \(e^t\) then subtract 1 then divide by t then integrate term by term

OpenStudy (anonymous):

you got that?

OpenStudy (anonymous):

\[ e^{t} = \sum_{n=0}^{\infty} t^{n}/n!\]

OpenStudy (anonymous):

I use this to expand e^t?

OpenStudy (anonymous):

\[e^{t}-1 = \sum_{n=0}^{\infty} t^{n}/n! -1\] so \[e^{t}-1/t = \(1/t)sum_{n=0}^{\infty} t^{n}/n! -1\]

OpenStudy (anonymous):

Then I integrate term by term ?

OpenStudy (anonymous):

\[e^t=1+t+\frac{t^2}{2!}+\frac{t^3}{3!}+...=\sum\frac{t^n}{n!}\]

OpenStudy (anonymous):

\[e^t-1=t+\frac{t^2}{2}+\frac{t^3}{3!}+..=\sum_{n=1}^{\infty}\frac{t^n}{n!}\]

OpenStudy (anonymous):

divide by \(t\) and get \[\frac{e^t-1}{t}=1+\frac{t}{2}+\frac{t^2}{3!}+...=\sum_{n=0}^{\infty}\frac{t^n}{(n+1)!}\]

OpenStudy (anonymous):

then integrate term by term

OpenStudy (anonymous):

\[t+\frac{t^2}{4}+\frac{t^3}{3\times 3!}+...\] make up a nice formula and you are done

OpenStudy (anonymous):

^ this guy = #winning

OpenStudy (anonymous):

I have to make up a formula from this or I'm done at this step?

OpenStudy (anonymous):

best make up a formula of course i don't mean "make up" i mean find it

OpenStudy (anonymous):

How do I do that :/